OpenStudy (anonymous):

let f(x)=(x^2+3x-4) and g(x)=(x+4). find f x g and state the domain. find f/g and state the do main.

OpenStudy (anonymous):

*domain

OpenStudy (austinl):

Hey there, welcome to OpenStudy. If you have any questions about the site feel free to shoot me a message. And when asking a question, you can edit the original blue area post. Now on to your question. \(f(x)=(x^2+3x-4)\) \(g(x)=(x+4)\) Part (a): \(f(x)\times g(x)\) You have any thoughts on how to set about this part?

OpenStudy (anonymous):

I know you use the the foil method to get the answer but the setting it up part is where im in need of help.

OpenStudy (austinl):

\(f(x)\times g(x)\Rightarrow(f\times g)(x)\) Just multiply the two of them together. Doing so requires quite a bit of distribution. You should end up with 6 terms when done.

OpenStudy (anonymous):

okay im lost... what is (fxg)(x).. how do I get that? like I tried multiplying the equations but its not what your saying